### v2

# DCA-Bench: A Benchmark for Dataset Curation Agents

Abstract.
The quality of datasets plays an increasingly crucial role in the research and development of modern artificial intelligence (AI). Despite the proliferation of open dataset platforms nowadays, data quality issues, such as incomplete documentation, inaccurate labels, ethical concerns, and outdated information, remain common in widely used datasets. Furthermore, these issues are often subtle and difficult to be detected by rule-based scripts, therefore requiring identification and verification by dataset users or maintainers–a process that is both time-consuming and prone to human mistakes. With the surging ability of large language models (LLM), it’s promising to streamline the discovery of hidden dataset issues with LLM agents. To achieve this, one significant challenge is enabling LLM agents to detect issues in the wild rather than simply fixing known ones. In this work, we establish a benchmark to measure LLM agent’s ability to tackle this challenge. We carefully curate 221 real-world test cases from eight popular dataset platforms and propose an automatic evaluation framework using GPT-4o. Our proposed framework shows strong empirical alignment with expert evaluations, validated through extensive comparisons with human annotations. Without any hints, most competitive Curator agent can only reveal $\sim$30of the data quality issues in the proposed dataset, highlighting the complexity of this task and indicating that applying LLM agents to real-world dataset curation still requires further in-depth exploration and innovation. The data and code are available at https://github.com/TRAIS-Lab/dca-bench.
Dataset Curation; Large Language Models; Automatic Evaluation
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737422isbn: 979-8-4007-1454-2/2025/08ccs: General and reference General conference proceedingsccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligenceccs: Information systems Data management systems
1. Introduction
High-quality datasets have become increasingly crucial for advancing artificial intelligence (AI) (Jain et al., 2020; Kaplan et al., 2020). Open dataset platforms, such as Hugging Face (Lhoest et al., 2021) and BIG-Bench (Srivastava et al., 2023), have substantially accelerated AI research and development by facilitating community contributions. However, community-contributed datasets often encounter subtle data quality issues, including insufficient documentation (Yang et al., 2024), inaccurate annotations (Klie et al., 2024), and ethical concerns (Gebru et al., 2021).
There have been existing efforts on standardizing dataset management practices (Gebru et al., 2021; Wang et al., 2023; Gan et al., 2024) and developing dataset curation toolkits and systems (Gupta et al., 2021; Mao et al., 2023). However, these toolkits and systems often rely heavily on rule-based scripts, which lack the necessary flexibility to detect the aforementioned subtle data quality issues. Consequently, existing techniques remain insufficient to address the complex challenge of automatically curating datasets at scale on open dataset platforms, where we still primarily rely on dataset users or platform maintainers as dataset curators to identify data quality issues.
Figure 1. Illustration of instances in DCA-Bench. The Issue Card displays a specific test case, including metadata such as Issue ID, source platform (e.g., BIG-Bench), type, associated tags, and hints. Relevant dataset files can be found using the Issue ID. Furthermore, DCA-Bench incorporates documentation from the dataset platform along with additional reference materials related to dataset curation and quality. We provide a convenient API to access data of each test case as well as reference materials. The Curator is asked to detect the issues in files by describing the issue context and pinpointing the location in the file where issues occur. The elements labeled title, content, and involved serve as ground truth for evaluating the Curator’s performance and are hidden from the Curator during testing.
Recent advancement of large language models (LLMs) has led to promising development of LLM agents for real-world software engineering problems (Jimenez et al., 2024a). For certain issues in software development, LLM agents have been shown to be capable of autonomously generating proper code to fix them (Mesh, 2024; Tao et al., 2024). Given this development, we envision that LLM agents could become an effective technique for building autonomous dataset curation systems.
However, there is still a significant gap between the need of dataset curation and the state-of-the-art LLM agents for software engineering. Specifically, existing studies primarily focus on developing agents that solve identified and well-defined issues, while dataset curation requires one to discover hidden issues in the community-contributed datasets—a distinct capability different from problem-solving (Jay, 1996; Wikipedia contributors, 2024) and remaining under-explored.
In response to these challenges, we introduce the Dataset Curation Agents Benchmark (DCA-Bench) as an initial step towards achieving autonomous dataset curation. Fig.1 illustrates the instance structure in DCA-Bench. Specifically, DCA-Bench aims to provide a comprehensive benchmark for evaluating LLM agents’ capability to discover data quality issues across online dataset platforms, the first step of the curation pipeline. Henceforth, we will consistently refer to such an LLM agent as a “Curator” to highlight its role in this task. A well-performed Curator can detect and locate existing issues, which is critical for a follow-up fix by human maintainers or other LLM agents.
Rather than defining the data quality issue beforehand, we adopt a bottom-up methodology for issue collections and classifications. Specifically, we collect 221 diverse and representative data quality issues from 8 online dataset platforms, covering a broad spectrum of problems such as data errors, documentation issues, file discrepancies, and legal/ethical risks. These issues are further categorized into 4 types and detailed with 18 tags, reflecting their varied content and complexity. The main features of DCA-Bench include:
  • Real-world Cases with Minimal Simplification: All test cases of DCA-Bench have references to real-world sources, allowing benchmark users to better understand them in the wild. To test the Curator’s ability under the complex real-world environment, each test case in DCA-Bench not only contains those flawed ones, but also files that without known issues.
  • Multiple Difficulty Levels: DCA-Bench provides four levels of hints for each test case in the benchmark. From a higher level hint, the Curator gains more information about the content and location of the issue. The motivation is to make the task more achievable and also gauge the information required for the Curator to detect these issues.
  • Accurate Automatic Evaluation: Unlike traditional machine learning tasks, the task of dataset curation does not have labels that can be directly evaluated by scripts. Human-level efforts are required to rate the performance of the Curator, which is not scalable. Therefore, we develop an automatic and accurate evaluation scheme using LLMs to replace human annotators.
In addition to the benchmark, we implement several straightforward baseline Curators for this task, including recent Open AI models (OpenAI, 2024a; OpenAI et al., 2024), Deepseek R1 (DeepSeek-AI et al., 2025a) and V3 (DeepSeek-AI et al., 2025b), and Meta Llama3.3-70B-Instruct (AI, 2024). The most competitive baselines only succeeds in detecting $\sim$ 29.86% issues without hints and 78.28% when given the most specific hint, demonstrating the difficulty of this benchmark.
We believe DCA-Bench will serve as a foundational initial step towards developing a fully autonomous and powerful dataset curation system, further enhancing the quality of community-contributed open datasets. This benchmark can also serve as a testbed for evaluating LLMs’ capability of problem discovery in addition to problem-solving, which is a critical area that has been under-explored.
We organize this paper as follows: In Section 5, we introduce the background of the dataset curation problem and existing literature. In Section 2, we go through an overview of the composition of DCA-Bench, the construction process, task definition, and the automatic evaluation framework. In Section 3, we carry out experiments to validate the reliability of our automatic evaluation scheme and discuss preliminary testing results of baseline Curators on DCA-Bench.
2. The Dataset Curation Agent Benchmark
In this section, we start with an overview of the proposed DCA-Bench in Section 2.1. Next, we introduce the benchmark construction process in Section 2.2 and give a detailed task definition in Section 2.3. Finally, Section 2.4 presents our evaluation framework.
2.1. Dataset Overview
The dataset assets provided in DCA-Bench consist of test cases and reference resources. Tab. 1 shows basic statistics of test cases. Each test case typically contains multiple files, designed to create a minimal environment for uncovering hidden data quality issues. We classify these test cases into four categories based on the number of files included and number of hidden issues in a test case. Additionally, we assign 18 descriptive tags to further label the cases, such as “data-problem”, “document-problem”, “infrastructure-problem”, and “ethical/legal-risk”. Furthermore, DCA-Bench provides reference resources, including related documentation and external materials such as dataset curation tutorials, to help users enhance Curator performance using methods like RAG (Lewis et al., 2021).
Table 1. Basic statistics of DCA-Bench. #Token is calculated using the tiktoken package to tokenize the file content. Non-text files are skipped. See Appendix E.2 for a detailed calculating procedure. For tag-level statistics, each test case can contain more than one tag. The definitions of types, tags and detailed statistics are provided in Appendix A.2, along with examples in Appendix A.2.
Statistic Number
Sample-Level #Samples $221$
Avg. #Files/Sample $2.13$
Avg. #Tokens/Sample $3.58\times 10^{6}$
Type-Level Single-Issue Single-File $61$
Single-Issue Multi-File $100$
Multi-Issue Single-File $14$
Multi-Issue Multi-File $46$
Tag-Level data-problem $197$
document-problem $83$
infrastructure-problem $19$
ethical/legal-risk $10$
Table 2. The number of tags in DCA-Bench. Note that the sum of the sub-category might not equal the parent category. For instance, if there’s an ethical or legal risk identified in a dataset document, it should be tagged as [“document-problem”, “ethical/legal-risk”], without including any sub-categories of ”document-problem” that are not relevant.
Category Number Sub-category Number
typo 18
wrong-format 14
inappropriate-file 4
ethical/legal-risk 10
internal-discrepancy 21
cross-file-discrepancy 44
data-problem 197 wrong-value 71
missing-value 15
data-leakage 2
apparent-corruption 40
hidden-corruption 59
document-problem 83 wrong-info 27
insufficient-info 52
infrastructure-problem 19 data-access 4
script-code 15
2.2. Test Case Construction
The test cases in DCA-Bench are collected from the real issues reported by users or maintainers of eight dataset platforms, including Hugging Face Dataset (Lhoest et al., 2021), BIG-Bench (Srivastava et al., 2023), Kaggle (Kaggle, [n. d.]), OpenML (Vanschoren et al., 2014), TensorFlow Dataset (Abadi et al., 2015), Open-Data-Registry (AWS-Labs, 2024), Five-Thirty-Eight (FiveThirtyEight, 2024), and Graph Learning Benchmark (GLI) (Ma et al., 2022). The construction of the test cases of the DCA-Bench involved the following two stages.
Issue Collection and Preprocessing
At this stage, we select and preprocess relevant and manageable data quality issues with the following steps:
  1. (1)
    Selection of Dataset Platforms and Issues: We focus on dataset platforms where users and maintainers can interact. For example, there is a “Discussion” section in each Kaggle dataset for discussing dataset-related issues, which helps collect meaningful data quality issues. We then manually collect data quality issues from these discussions. Please refer to A.3 for more details about the choice of platforms and case selections.
  2. (2)
    Issue Filtering and Modification: We prioritize the issues that have received feedback from dataset maintainers. Otherwise, we manually verify the problems mentioned in the issues to ensure their validity. We exclude issues that are not directly related to dataset files uploaded by contributors (e.g., issues about Tensorflow or Hugging Face data loading APIs). For issues involving multiple sub-issues or file sizes over 512MB, we either discard them or split them into manageable sub-cases with adjustments.
  3. (3)
    File Downloading: After gathering candidate issues, we download the dataset files from the version before the problems are fixed. In order to simulate real-world scenarios, the dataset files included in each test case are not limited to those involving issues.
Hints Generation
Asking Curators to discover hidden issues from raw dataset files can be very challenging. To make it more manageable and test the Curators’ capability in finer granularity, we generate different levels of hints to help the Curators locate the hidden issues. We apply GPT-4 with carefully designed prompts [^1] See Appendix B.8 for prompts used for the Hint Writer. to generate three levels of hints for each test case, in addition to the no-hint setting:
  • $h_{0}$: No hint provided. In this case, the Curator is required to detect the issue fully on its own.
  • $h_{1}$: General description of the issue, without any specific details or hints on the location.
  • $h_{2}$: Information about which files are involved in the issue, in addition to information from $h_{1}$.
  • $h_{3}$: Partial contextual information about the issue, in addition to information from $h_{2}$.
After generating the hints, we manually double-check the hints and make necessary modifications to ensure that the generated hints follow the guidelines above.
2.3. Benchmark API
After gathering test cases and related information from dataset platforms, as well as generating multi-level hints, we proceed to develop the benchmark API, which encapsulates the inputs to the Curator and the Evaluator, adhering to the Task I/O paradigm as illustrated in Fig.2.
In this framework, the Curator receives following from benchmark API: (i) dataset files with hidden issues, (ii) hints on the issues, and (iii) optional dataset documentation and reference materials. It is then asked to identify, describe, and provide contextual evidence for these issues. The Curator is allowed to use any strategy and tools to process provided dataset files and find the issue, e.g., flattening the file contents and feeding them to the Curator, applying RAG technology, using a code interpreter to write and execute programs, or combining multiple strategies together.
The Evaluator then compares the outputs from the Curator with the label provided by the benchmark API to rate the performance of the Curator. The performance of the Curator is then classified into three levels: fail, success, and success+, which we introduce in the following section.
Figure 2. The Task I/O of DCA-Bench. For each test case, the input for the Curator includes dataset files and hints, with reference materials and platform documentation being optional. The Curator is then required to provide a description of the issue and corresponding contextual evidence. The label of the test case includes the issue title, content, the involved file names, and corresponding contextual evidence. Given the output from the Curator and the label, the Evaluator is then asked to rate the performance of the Curator.
2.4. Evaluation Pipeline
We now discuss the motivation and design of our evaluation pipeline. We evaluate the Curators in terms of the following two aspects: (i) whether the Curators accurately identify the issues, and (ii) whether the Curators provide necessary contextual evidence about the issue. We accordingly categorize the performance of the Curator into three levels:
  • fail: The Curator fails to discover any issues, only identifies irrelevant issues, or acknowledges the issue but offers completely wrong contextual evidence.
  • success: The Curator identifies the annotated issue and provides at least one correct piece of contextual evidence.
  • success+: The Curator correctly identifies all issues and provides all necessary contextual evidence.
Note that this evaluation process has several intrinsic challenges. Firstly, due to the nuanced nature of text generation, different Curator outputs can refer to the same issue, while similar Curator outputs may also point to different issues. Fixed keyword (He et al., 2024a) or rule-based code tests (Jimenez et al., 2024a) are often insufficient to distinguish the nuances. Additionally, Curators may be able to identify the issues using only a portion of the contextual evidence annotated in the test case, which means that evaluation protocols requiring the inclusion of all the annotated contextual evidence can be overly stringent. While human experts can easily capture such nuances and properly evaluate the Curator performance, human evaluation is too expensive for a benchmark in practice. Recognizing that evaluating the Curator’s performance is easier than building the Curator agent, we propose an automatic evaluation pipeline where LLMs are leveraged to serve as the Evaluator.
We explored several prompting strategies using a few samples of human-annotated outcomes on a handful of issues and ended up with the following prompt design, which asks the LLMs to rate the Curator outputs in terms of a few criteria, and then aggregate these ratings to obtain an overall score for the Curator performance. The criteria we have in the Evaluator prompt are listed as follows:
  • Precise Contextual Evidence: This criterion evaluates whether the Curator has detected and described the issue while providing accurate contextual evidence. The Curator should receive a low rating if it identifies the issue but fails to locate the correct context. A medium or high rating is justified only if the Curator identifies the issue and accurately pinpoints where it arises.
  • Detailed Issue Analysis: This criterion examines whether the Curator’s response extends beyond the mere repetition of provided hints, indicating a profound understanding of the issue.
  • Relevance of Reasoning: This criterion ensures the Curator’s logical reasoning is directly pertinent to the addressed problem rather than being generic or irrelevant.
For each test case, the Evaluator receives the Curator’s output along with the hints and the ground truth annotations about the issues in the test case. Then, the Evaluator is asked to rate the Curator’s performance for each criterion with a real number from 0 to 1. The final score is a weighted sum of the three ratings, with weights of $0.85$, $0.15$, and $0.05$, reflecting their respective importance. The Evaluator then categorizes the Curator’s performance according to defined thresholds on the final score [^2] More details about the settings of weights and thresholds are provided in Appendix B.1..
While it’s natural to ask LLMs to directly classify the Curators’ performance into fail, success, success+, this strategy is poorly aligned with the annotation of human experts as we tested. In addition to common prompt methods (e.g., COT (Wei et al., 2023), role-playing (Kong et al., 2024), delimiter (OpenAI, 2024c)), we believe more prompt engineering is needed to improve the performance of the Evaluator, and here is our design:
We chose to have the Evaluator rate on specific metrics because this strategy allows for iterative refinement of the prompt by adjusting the weight and threshold of each metric. This approach decreases uncertainty by introducing fixed numerical comparisons, which provides a clearer direction for tuning, as opposed to relying solely on textual standards, which can be more ambiguous.
In designing the Evaluator’s prompt, we started by selecting a small subset of human-labeled samples. We began with an initial weight distribution of ($0.7,0.2,0.1$) to reflect the relative importance of each metric in Section 2.4. The high initial weight for ”Precise Contextual Evidence” was chosen to align with our focus on whether the Curator identifies annotated issues by providing contextual evidence, as discussed in Section 2.4. We then fine-tuned these weights using the selected subset to better align the Evaluator’s prompt with human judgments.
The motivation behind this design is to make iterative refinement of prompts more effective. Developing a good prompt requires testing it on a subset of human-labeled issues with corresponding outcomes from Curators and then making adjustments based on the outcomes. Unlike the model training process, which benefits from clear gradients and systematic optimization, tuning text prompts is often an empirical process necessitating numerous trials without the guarantee of improvement in its performance. Employing numerical weights and thresholds could enable us to tune the prompt more fine-grained, only changing the weights or threshold based on the results, which share similarities with linear regression, resulting in a more efficient prompt tuning process and better performance. By ”training” on a this human-labeled subset, we finalized weights in the prompt as $0.85$, $0.15$, and $0.05$.
To enhance the accuracy of the evaluation, a simplified voting strategy (Verga et al., 2024) was implemented to mitigate the influence of randomness. The strategy comprises three rounds of voting. In the first round, we collect results from $n$ Evaluators. If a consensus is reached, the decision is made with the consensus; however, if there is no consensus, we conduct another round of voting within $m$ Evaluators, and apply a majority vote to make the decision. In the event of a tie, additional votes will be conducted one by one until a definitive majority vote is determined. In practice, $m$ and $n$ are both set to 2.
The testing results of the Evaluator on larger scale samples are displayed in Section 3.1, demonstrating its reliable performance. Notably, we collect annotated test data for the Evaluator after we have finalized the prompt design for our Evaluator. Therefore, we believe the proposed prompt design of our Evaluator can generalize across different Curators and issues in the DCA-Bench.
3. Experiments
In this section, we first assess the performance of the Evaluator, demonstrating its strong alignment with human annotations, along with its robustness and negligible bias. We then utilize this Evaluator to evaluate baseline Curators on DCA-Bench and further analyze the results.
3.1. Validation of the Evaluator
3.1.1. High Alignment with Human
To verify the effectiveness of the proposed agent Evaluator, we conduct a comparative analysis against human annotations. We randomly select 23 test cases for each of the four hint levels, resulting in a total of 92 (issue, hint) pairs. Using these pairs, we collect the outputs of a baseline Curator agent, which is based on the OpenAI Assistant API with GPT-4-0125-preview equipped with the Code Interpreter tool [^3] See https://platform.openai.com/docs/assistants/tools/code-interpreter., on each of the 92 pairs. The Evaluator and human annotators then rate the Curator outputs independently. To develop a reliable and cost-efficient Evaluator, we test following different models as backends: GPT-4o-2024-11-20, GPT-4o-2025-05-13 (OpenAI et al., 2024), o3-mini-2025-01-31 (OpenAI, 2024b), GPT-4-0125-preview (OpenAI, 2024a), GPT-3.5-Turbo (OpenAI, 2023a), Llama3.3-70B-Instruct (AI, 2024), Llama3-70B-Instruct(Dubey et al., 2024), Deepseek-R1 (DeepSeek-AI et al., 2025a), and Deepseek-V3 (DeepSeek-AI et al., 2025b). Details for setting up those Evaluators are included in Appendix B.2. Human annotators follow a structured codebook [^4] Please see Appendix B.4 for details to ensure rating consistency. Finally, we compare the Evaluator’s ratings with human annotations (ground truth) using a binary classification scheme: fail versus succ (where succ combines both success and success+). Additional results based on a three-class classification scheme (fail, success, and success+) are provided in Appendix B.5.
Table 3. The performance of Evaluators with different backend models, using human annotations as ground truth. We treat succ as the positive class when calculating Precision, Recall, and F1 Score. The “$\kappa$ Value” refers to Cohen’s $\kappa$ between the Evaluator ratings and the human annotations. Color Legend:   (95–100%),   (85–95%),   (60–85%),   (0–60%). Based on the table, gpt-4o-2024-11-20 and gpt-4-0125-preview outperform other models. Among open-source models, DeepSeek R1 and DeepSeek-V3 demonstrate reasonably good performance, though there remains a noticeable gap compared to the top-performing models. Conversely, LLama-3-70B-Instruct and GPT-3.5-Turbo exhibit the weakest performance. Given that gpt-4o-2024-11-20 is more cost-efficient than gpt-4-0125-preview, we primarily adopt it for evaluating Curators in Section 3.2.
Success Rate / %
Model Name Accuracy Precision Recall F1 Score $\kappa$ Value
gpt-4o-2024-11-20 97.83 100.00 92.59 96.15 94.64
gpt-4-0125-preview 96.74 92.86 96.30 94.55 92.22
gpt-4o-2024-05-13 92.39 81.25 96.30 88.14 82.59
DeepSeek-R1 92.39 81.25 96.30 88.14 82.59
DeepSeek-V3 93.48 88.89 88.89 88.89 84.27
gpt-3.5-turbo 68.48 48.21 100.00 65.06 42.15
Meta-Llama-3-70B-Instruct 69.57 49.09 100.00 65.85 43.68
Meta-Llama-3.3-70B-Instruct 88.04 72.22 96.30 82.54 73.73
o3-mini-2025-01-31 91.30 100.00 70.37 82.61 77.04
As shown in Tab. 3, Evaluators using all backend models achieve high recall scores, indicating that they rarely misclassify succ as fail. However, weaker models such as GPT-3.5-Turbo and Llama-3-70B-Instruct frequently misclassify fail as succ, as reflected in their relatively low precision scores.
Overall, Evaluators with the GPT-4o-2024-11-20 and GPT-4-0125-preview backends exhibit strong alignment with human annotations, suggesting that they can serve as reliable proxies for human evaluations in our benchmark. Other models, such as DeepSeek-R1 and DeepSeek-V3, demonstrate good performance, though a noticeable gap remains compared to the top two models.
3.1.2. Consistent Evaluation
To assess the consistency of ratings generated by our prompting and voting strategy, we conducted three independent evaluations using two Evaluators powered by GPT-4-0125-preview and GPT-4o-2024-11-20. Each Evaluator rated the same outputs produced by the GPT-4-0125-preview Curator on a subset of DCA-Bench (92 data-hint combinations). The results show high stability, with a standard deviation of ±$\pm 0.89\%$ for GPT-4-0125-preview and ±$\pm 0.51\%$ for GPT-4o-2024-11-20. Detailed numerical results are provided in Tab. 7 in the Appendix.
3.1.3. Negligible Self-Preference and Length Bias
To ensure the fairness of our evaluation strategy, we further examine two common biases that often arise when using LLMs as evaluators: self-preference bias and length bias. Through a thorough analysis, we find that our evaluation strategy remains robust against both biases.
Self-Preference Bias: We first assess under such designs, whether the Evaluator favors Curators that use the same model as their backend. Using GPT-3.5-Turbo and GPT-4-0125-preview as both Curators and Evaluators, we compute the Average Odds Difference (AOD) (Majumder et al., 2022), a fairness metric measuring disparities in true and false positive rates. The results indicate an AOD of $6.09\,(\pm 1.72)\%$ for GPT-3.5-Turbo and $5.31\,(\pm 2.76)\%$ for GPT-4-0125-preview, both well within the fairness threshold of $10\%$. Furthermore, GPT-4-0125-preview Evaluator aligns closely with human annotations across multiple evaluation metrics, confirming the negligible impact of self-preference bias.
Length Bias: To test whether response length influences evaluation outcomes, we modify Curator responses to be more verbose or succinct while preserving content. Experiments using GPT-3.5-Turbo-0125 and GPT-4o-mini-0718 show minimal differences in accuracy across different lengths (at most $\approx 1\%$ variation), indicating that response length does not significantly affect evaluations.
These findings confirm that our LLM Evaluator maintains fairness with negligible self-preference and length bias. Please see more details in Appendix B.7.
3.2. Benchmarking the Baseline Curators
3.2.1. Experimental Setup
In the following sections, we use the Evaluator powered by GPT-4o-2024-1120 due to its lower cost and strong performance, as shown in Tab.3. We also implement several baseline methods with different backend models, detailed in Tab.4.
Table 4. The performance of several Curators with different backend models across various hint levels on the full DCA-Bench dataset is evaluated. The Evaluator is based on the GPT-4o-2024-11-20 model. The success rate here refers to succ (success and success+, defined in Sec. 2.4). Notably, without any hints ($h_{0}$), even leading models such as DeepSeek-R1 and GPT-4o-2024-11-20 perform unsatisfactorily, detecting only about 30% of issues. As more informative hints are provided, the success rates increase accordingly. However, even with the most informative hints, the success rate of the models fails to exceed 80%. The results also indicate that incorporating RAG with knowledge relevant to dataset curation does not guarantee a performance improvement across all test cases and models.
Success Rate / %
Model Name $h_{0}$ $h_{1}$ $h_{2}$ $h_{3}$ Avg.
w/o Knowledge RAG
DeepSeek-R1 29.86 52.04 52.04 73.30 51.81
DeepSeek-V3 15.84 39.82 39.82 76.02 42.87
GPT-4-0125-preview 10.86 27.15 34.84 67.42 35.07
GPT-4o-2024-11-20 20.36 41.63 47.51 76.02 46.38
w/ Knowledge RAG
DeepSeek-R1 29.41 45.25 56.11 78.28 52.26
DeepSeek-V3 12.22 38.91 42.99 75.57 42.42
For open-source models, we utilize SGLang (Zheng et al., 2024) for local deployment and LangChain (Chase, 2022) to manage file-based context inputs. One intuitive way to improve the Curator’s performance is by incorporating external knowledge about dataset curation. To achieve this, we implement a RAG pipeline using LangChain. The RAG vector store is built using materials from DCA-Bench, as described in Section 2.1. Among commercial models, we access GPT-4-0125-preview via the OpenAI Assistant API, which includes the Code Interpreter tool. [^5] As noted by OpenAI API users, the performance of the OpenAI Assistant API can fluctuate. See Appendix E.1 for further details. For GPT-4o-2024-1120, we use the LangChain interface in the same way as for open-source models. Additional implementation details are provided in the Appendix B.3.
3.2.2. Results and Analysis
Tab. 4 presents the results of various baseline Curators evaluated using the Evaluator powered by the GPT-4o-2024-11-20 model. Among them, the Curator utilizing DeepSeek-R1 with RAG enabled achieves the highest average score of $53.05\%$, while GPT-4-0125-preview accessed via the OpenAI Assistant API performs the worst. Notably, without any hints ($h_{0}$), even the strongest models, such as DeepSeek-R1 and GPT-4o-2024-11-20, struggle to detect issues, achieving a success rate of only around 30%. As more informative hints are introduced, the success rates improve accordingly. However, even with the most informative hints, no model surpasses an 80% success rate.
We also observed that, although using RAG improves the performance of DeepSeek-R1, the improvement is not significant. Moreover, when applying the same RAG approach to the DeepSeek-V3 model, we even observed a decline in performance. Therefore, while leveraging RAG with dataset curation materials that provide curation knowledge is an intuitive approach, we believe it will need more designs and exploration to offer practical assistance to LLM agents in solving real-world dataset challenges. This could be attributed to several factors. We suspect that the reference materials introduce overly generic information, which occupies the model’s context window and diverts its attention from the specific issues at hand. Additionally, the diversity of these materials may not be sufficient to fully cover the range of issues present in DCA-Bench.
Thus, we believe that integrating more extensive and specialized knowledge, leveraging advanced RAG techniques, and fine-tuning models have great potential for improvement. Research on multi-agent collaborations (Wu et al., 2023a) also shows promising prospects. Further exploration is needed, and we leave it for future works.
Besides, we analyze the success rate of the Baseline Curator on issues with different token lengths. We observe that the success rate doesn’t monotonically decrease as content length grows. A detailed analysis of its performance across different issue types and hint levels is provided in Appendix C.
Overall, the results of the baseline Curators suggest that the proposed DCA-Bench benchmark presents a substantial and meaningful challenge for the AI community, highlighting the need for further exploration and research.
4. Ethics Statement
Our study involves human annotators, who are the authors themselves. The datasets utilized in DCA-Bench are sourced from various platforms, with proper licensing documentation provided. Additionally, we have compiled a detailed table (DCA_Benchmark.csv) in the supplementary materials, listing license information for all data files. Regarding secondary usage of user-generated data, we reviewed the terms of service for GitHub, HuggingFace, and Kaggle. Based on their policies and existing academic precedents, we believe our usage aligns with acceptable practices for non-commercial research purposes. We do not collect any user information and commit to removing content upon valid request.
Furthermore, our dataset includes 10 instances involving ethical or legal risks, with 2 data points potentially exhibiting bias. These cases are included to assess LLMs’ ability to identify and address ethical concerns, contributing to a realistic evaluation of AI safety. The annotation process poses minimal risk to annotators. Please see more details regarding the ethical considerations in Appendix D.
5. Related Work
Dataset Quality Management. With the advancement of modern AI, the need for high-quality datasets is surging (Jain et al., 2020; Kaplan et al., 2020). Consequently, open dataset platforms like Hugging Face and Kaggle keep growing rapidly. However, recent studies (Klie et al., 2024, 2023; Weber-Genzel et al., 2023) indicate that many popular datasets suffer from errors, biases, or annotation artifacts (Gururangan et al., 2018), and lack proper documentation (Yang et al., 2024). In addition, common issues in real-world datasets include problems with data loading scripts and file discrepancies, where conflicting information exists between files.
There have been many existing efforts on dataset quality management. The FAIR principles (Wilkinson et al., 2016) provide a broad framework for creating and managing datasets. Gebru et al. (2021) puts forward a standardized process for documenting datasets in the context of machine learning. Gong et al. (2023) propose several strategies, including dataset profiling, data cleansing, and quality monitoring. More recent works (Gebru et al., 2021; Wang et al., 2023; Gan et al., 2024) further explore standardizing dataset management practices. Based on these works, researchers have developed advanced dataset management toolkits and systems (Gupta et al., 2021; Mao et al., 2023; Zhou et al., 2024). More recently, Zhou et al. (2023); Chen et al. (2023b) developed systems that are capable of greatly enhancing data quality. However, they focus on filtering a large-scale LLM-pertaining corpus rather than detecting dataset quality issues. There are some specialized works focused on detecting harmful content (Althobaiti, 2022; Kirk et al., 2023), annotation errors (Weber-Genzel et al., 2023; Wang and Mueller, 2022) and data attributions (Longpre et al., 2023). Nevertheless, those methods are not generalizable enough and lack the flexibility to solve challenges such as nuanced ethical biases (see example A.1) and incorrect annotations involving logical or factual errors (see example A.3). Moreover, studies on file discrepancies (example A.2) and relevant issues remain scarce. Consequently, we still primarily rely on dataset users or platform maintainers to identify these issues in practice. We envision LLM agents as a promising technique to address the challenges in this space and propose a benchmark dataset as an initial step toward this goal.
LLMs for Software Engineering. Our work is also relevant to the emerging research direction on using LLMs to solve software engineering tasks. Fan et al. (2023) suggest that there are two major categories in this area: (i) code generation and completion, and (ii) maintenance and evolution.
There has been extensive literature on LLMs for code generation and completion, in terms of both models and benchmarks. On the model side, various code generation LLMs have been proposed (Chen et al., 2022; Luo et al., 2023; Rozière et al., 2024), along with agent frameworks designed to tackle complex coding problems, such as data analysis (Hong et al., 2024; Guo et al., 2024). On the benchmark side, datasets have been created to complete functions and short programs from natural language descriptions (Jimenez et al., 2024a; Chen et al., 2021; Austin et al., 2021; Lai et al., 2022; Zhang et al., 2024; Ding et al., 2023) and retrieve relevant code (Liu et al., 2023b). On the framework side, Reflexion (Shinn et al., 2023), Parsel (Zelikman et al., 2023) and MetaGPT (Hong et al., 2023) are the best reasoning frameworks at the time of writing. We note that the main purpose of the proposed DCA-Bench is to measure LLM agents’ capability of discovering hidden dataset issues instead of generating code, although generating and running some test code may be helpful.
For the second category, maintenance and evolution, it involves tasks such as localizing and fixing bugs (Wu et al., 2023b; Feng and Chen, 2023), improving program performance (Garg et al., 2024; Chen et al., 2023a), and refactoring code without changing program behavior (Poldrack et al., 2023). Our work falls into this category while differing from previous studies in two key aspects. Firstly, while most literature focuses on tasks with well-defined outcomes, such as fixing a buggy function, our work addresses non-standard issues like ethical concerns in data points or cross-file inconsistencies. Secondly, few studies have explored using LLMs to maintain dataset repositories, which is increasingly important for the machine learning community.
LLMs as A Proxy of Human Evaluation. Recent studies have shown that LLMs, when carefully prompted, can serve as a good proxy of human evaluations for a number of scenarios, such as evaluating text generation quality (Liu et al., 2023a; Chiang and Lee, 2023), reasoning ability (He et al., 2024b; Hao et al., 2024), and generated image quality (Chen et al., 2023c; You et al., 2024). It’s difficult to automatically evaluate the agents’ performance in complex generative tasks like text generation. Traditional reference-based metrics, such as BLEU and ROUGE, have been shown to have a relatively low correlation with human judgments (Liu et al., 2023a). These problems necessitate human intervention in evaluation judgment, making the process costly and hard to scale up. Facing these challenges, researchers have explored using LLMs for evaluations. Several existing works have shown that well-designed LLM evaluations align well with human evaluations through techniques including auto COT (Chiang and Lee, 2023), self-explaining rating (Chiang and yi Lee, 2023), and voting with multiple LLM evaluators (Verga et al., 2024). More and more researchers have applied LLMs in the evaluation process, including evaluating text generation quality (Liu et al., 2023a; Chiang and Lee, 2023), reasoning ability (He et al., 2024b; Hao et al., 2024), and generated image quality (Chen et al., 2023c; You et al., 2024).
Inspired by these studies, we adopt an LLM to automatically evaluate the Curator’s responses based on carefully designed instruction. We also empirically show that our LLM-based Evaluator highly aligns with human preference, ensuring a reliable automatic evaluation scheme.
6. Conclusion
To help the development of LLM agents capable of dataset curation, we present DCA-Bench, a collection of representative data quality issue cases from popular dataset platforms. Instead of fixing predefined issues, DCA-Bench aims to test the agent’s capability to discover hidden data quality issues, a critical initial step in the dataset curation pipeline. To efficiently and effectively evaluate the performance of the Curator agents, we develop an LLM-based Evaluator with carefully designed prompts, which aligns well with human annotators. We conduct a benchmark study on the baseline Curator using DCA-Bench, and the results indicate that while current state-of-the-art LLMs have great potential in real-world dataset curation, further exploration and innovation are needed to fully realize their capabilities.
Limitations and Future Works. Dataset curation is a complex and comprehensive problem, and test cases we collect might not fully cover the entire problem set. Additionally, due to the complexity of dataset files, we cannot guarantee that all the issues in the dataset files in DCA-Bench are labeled. Lastly, we have not yet explored other modalities such as images or audio, which could enhance the curation of multimedia datasets. Future research could focus on developing more sophisticated LLM agent systems, establishing best practices for managing multi-modal information in dataset curation, or creating realistic simulation environments for testing LLM agents. We hope our work inspires these future studies, and we decide to leave these aspects for further works.
Appendix A Details about examples, statistics, and construction process of DCA-Bench
A.1. Real World Dataset Curation Examples
The real-world dataset curation process can be complex. During the dataset contribution procedure, authors and administrators must communicate iteratively to address issues in the contributed datasets. These issues can be numerous and challenging to address comprehensively at once, even for proficient individuals. Additionally, new issues often arise while fixing existing ones, leading to a complex contribution procedure. Below, we provide a few URLs to real-world cases on the TensorFlow dataset and BIG-Bench to illustrate these challenges:
Besides, dataset issues are often reported by dataset users, even though uploaded datasets have passed the initial checks of the administrators. Here are some representative examples classified by tags:
A.2. Issue Type and Tag
We classify the issues into single-issue / multi-files and single-file/multi-files.
  • single-issue: This sample has only one known issue.
  • multi-issue: this sample has multiple issues. If many issues of the same type exist in a sample, we also classify them into multi-issue.
  • single-file: the environment of this sample has only one file
  • multi-file: the environment of this sample has multiple files
It’s important to distinguish between the total number of files in a test case and the number of files affected by issues. Tab. 5 shows the distribution of issue-affected files in DCA-Bench.
Table 5. Statistics of files involved with issues. On average, each instance has 1.44 files involved in the issue.
Number of Issue-Affected Files Number of Test Cases
1 159
2 54
3 6
$>$ 3 2
To help benchmark users better understand the issue, we further assigned each of them several tags. Some tags have a structure of affiliations.
  • typo:  Issues caused by typographical errors.
  • wrong-format:  Problems related to incorrect formatting.
  • inappropriate-file:  Missing, empty or redundant files.
  • ethical/legal-risk:  Ethical or legal risks associated with the dataset, such as racial bias, missing license.
  • cross-file-discrepancy:  Discrepancies between files (e.g., meta-information in the documentation does not match the actual data).
  • internal-discrepancy:  Discrepancies within a single file.
  • data-problem:  Issues related to the data files.
      wrong-value:  Errors in the data values.
      missing-value:  there are missing columns or values.
      data-leakage:  Risks of data leakage.
      apparent-corruption:  Errors in the data files that can be easily detected using the information within the file. (e.g. duplicated data, apparently wrong target format compared to other targets in the same file, CSV file with wrong format)
      hidden-corruption:  Errors in the data files that require external knowledge and logical reasoning to resolve.
  • document-problem:  Issues related to the documentation files(e.g. README, DataCard, meta-data and other descriptive text) of the dataset.
      wrong-info:  Wrong information in the documentation files (e.g., wrong meta-information, typos, invalid URLs, and email addresses).
      insufficient-info:  Missing or unclear information in the documentation files (e.g., meanings of columns, labels, units of data values).
  • infrastructure-problem:  Issues with fetching, loading, processing, or displaying data.
      data-access:  Problems accessing the data.
      script-code:  Issues with scripts.
Example A.1: An issue example reported on Kaggle which involves racial bias
TitleBoston House Prices B feature is RACIST
Meta-Info
ID: 7e8f31cb-8c2a-4676-b3d4-941a64184a26 Platform: Kaggle Issue Type: single-issue & multi-file Issue Tags: ethical-legal-risk document-problem Source: https://www.kaggle.com/datasets/vikrishnan/boston-house-prices/discussion/429030
Content
B: 1000(Bk-0.63)2 where Bk is the proportion of blacks by town No other race is featured in this dataset. Red-lining anyone?
Involved Files - name: datacard.md - context: PTRATIO:pupil-teacher ratio by town 12. B: 1000(Bk-0.63)2 where Bk is the proportion of blacks by town 13. LSTAT:lower status of the population
Hints $h_{1}$ dataset contains potentially biased feature $h_{2}$ bias in a feature documented in a markdown file $h_{3}$ a feature in datacard.md is described using a formula that appears to single out one race
Example A.2: An issue example reported on BIG-Bench that involves a discrepancy between dataset files.
TitleMiss aligned static information
Meta-Info
ID: 80d6db6a-6cbf-4261-8d13-3244e7fb54fd Platform: BIG-Bench Issue Type: single-issue & multi-file Issue Tags: cross-file-discrepancy document-problem/wrong-info Source: https://github.com/google/BIG-bench/pull/498
Content
The stastic info in README.md is not aligned with the actual data file. There are 190 stories rather than 194 stories; 99 Ÿesr̈ather than 100 Ÿes;̈ 91 N̈or̈ather than 94 N̈o.̈
Involved Files 1. name: task.json - context: the number of datapoints in data files. 2. name: README.md - context: We collected 194 stories from 30 papers published in the span of 1989 to 2021. Each story has a causal judgment question associated with it with a "Yes" or "No" answer. We carefully balanced the dataset -- there are 100 "Yes" answers (52%) and 94 "No" answers (48%). Each paper that we collected from has conducted rigorous human experiments. We follow a simple binarization strategy to reflect the majority of human agreement and use it as the ground truth to evaluate the AI model.
Hints $h_{1}$ Mismatched quantitative information $h_{2}$ Mismatched quantitative information in README.md and data file $h_{3}$ Quantitative discrepancies between README.md and task.json, specifically in the counts of stories, ’Yes’ answers, and ’No’ answers
Example A.3: An issue example which has a wrong target label that needs precise factual knowledge to discern
TitleError in 118th Congress data
Meta-Info
ID: 51e12546-8bf3-473c-9ed6-f85d63c357ce Platform: FiveThirtyEight Issue Type: single-issue & multi-file Issue Tags: data-problem/hidden-corruption, data-problem/wrong-value Source: https://github.com/fivethirtyeight/data/issues/336
Content
The "congress-demographics" data includes Benjamin Eric Sasse as being a member of the 118th Congress but he resigned after the 117th.
Involved Files - name: data_aging_congress.csv - context: The "congress-demographics" data includes Benjamin Eric Sasse as being a member of the 118th Congress but he resigned after the 117th.
Hints $h_{1}$ inaccurate data entry $h_{2}$ an inaccurate data entry in a CSV file $h_{3}$ an entry in ’data_aging_congress.csv’ inaccurately includes a member as part of the 118th Congress
Example A.4: An issue example which has a wrong target label that requires translation ability to discern
TitleMistranslation in conlang_translation task?
Meta-Info
ID: bb29a2c3-872b-41cc-ac55-b26f22043da6 Platform: BIG-Bench Issue Type: single-issue & multi-file Issue Tags: data-problem/wrong-value, data-problem/hidden-corruption Source: https://github.com/google/BIG-bench/issues/553
Content
On the English to Gornam translation, one of the translations may be incorrect. Specifically, the example given is English: They want to eat my pizzas. Gornam: Sa wott min Pizzas atten. However, from looking at the other examples, it seems like when the subject is plural (like they), the suffix en is attached to the word, so the correct Gornam translation should be Sa wotten min Pizzas atten. Is this true, or is there some flaw in this logic that we are missing?
Involved Files - name: task.json - context: "examples": [ {"input": "I want to buy the orange.", "target": "Ek wott dei Orange leuren."}, {"input": "He is wearing my pants.", "target": "Ha trugt min Rose."}, - {"input": "They want to eat my pizzas.", "target": "Sa wott min Pizzas atten."}, + {"input": "They want to eat my pizzas.", "target": "Sa wotten min Pizzas atten."}, {"input": "I can eat.", "target": "Ek conn atten."}, {"input": "They eat their shirts.", "target": "Sa atten hir Pemts."}, {"input": "He tries on my coat.", "target": "Ha roturt min Enzu en."},
Hints $h_{1}$ incorrect translation example $h_{2}$ incorrect translation example in a JSON file $h_{3}$ a potential translation inconsistency in ’task.json’ involving plural subject examples
A.3. Discussion on Choices of Dataset Platforms and Cases Selections
We mainly focus on two types of dataset platforms. The first type includes HuggingFace and Kaggle, which hosts a huge volume of datasets and offers a dedicated discussion section for individual datasets. The second type is hosted on GitHub, such as BIG-Bench, where the data quality issues are identified from the issue or pull request (PR) sections. Roughly speaking, for HuggingFace and Kaggle, we first select a few popular datasets and then examine the discussion posts relevant to data quality issues. For GitHub-based repositories, we obtain data quality issues by browsing the issue or PR sections.
Our detailed collection process is outlined as follows:
Kaggle:
  1. (1)
    We begin by using Kaggle API to select datasets that have $\geq 5$ discussions with a proper files size $\leq 512$ MB, and collect them into a candidate list.
  2. (2)
    We manually review the discussion sections of these datasets, collecting discussions relevant to dataset issues and compiling them into a table.
  3. (3)
    For each collected discussion, if the dataset issues are confirmed by either the dataset owner or other participants (with at least two reports of the issue), we consider it as verified. If not, we manually verify the issues by checking the corresponding files, which is straightforward due to the dataset’s version history control (in case the issue has been resolved in the current version).
HuggingFace:
  1. (1)
    We follow a similar approach to Kaggle, first selecting candidate datasets that have enough discussions and proper file size by using HuggingFace API
  2. (2)
    Using the same strategy as for Kaggle, we collect discussions relevant to dataset issues and verify them through the same process.
Platforms Hosted on GitHub:
On GitHub, dataset issues are identified from several sources,
  1. (1)
    Issues raised by platform maintainers during the review of new dataset contribution PRs. These are inherently verified by the maintainers.
  2. (2)
    Issues reported by users in the issues section. If these issues are confirmed by the maintainers, they are considered verified (And it’s usually linked to a PR so as to fix it). If not, we manually check the corresponding files ourselves.
We have also surveyed on other dataset hosting platforms, which we found are not ideal for constructing DCA-Bench. Zenodo ([n. d.]) is a platform for safely storing and sharing research data, which fully follows FAIR principles (Wilkinson et al., 2016). However, to our knowledge, it doesn’t support a feedback mechanism, which makes it unclear whether there are issues within those dataset files and whether the descriptions align with the datasets. Compared to HuggingFace and Kaggle, this lack of feedback makes the platform less informative about the issues that concern dataset users, which is why we did not collect datasets from it. Harvard-Dataverse ([n. d.]) faces similar issues, as it has a GitHub repository but does not use it for reporting dataset problems. Domain-specific datasets like the Materials-Data-Facility ([n. d.]) or PANGAEA ([n. d.]) also share the same problem, making it difficult for us to collect issue samples from them, which is why they were less interesting for us. Open-Science-Framework ([n. d.]) seems not to be a standard dataset hosting platform—it is a platform for hosting projects, where the files do not necessarily have to be datasets, so we dismissed further discussions.
Appendix B Details about DCA-Bench Evaluator, Hints Writer and Curator
B.1. Design Details of the Evaluator
In this part, we elaborate on the prompt design of our Evaluator. While it’s natural to ask LLMs to directly classify the Curators’ performance into fail, success, success+, this strategy is poorly aligned with the annotation of human experts as we tested. In addition to common prompt methods (e.g., COT (Wei et al., 2023), role-playing (Kong et al., 2024), delimiter (OpenAI, 2024c), we believe more prompt engineering is needed to improve the performance of the Evaluator, and here is our design:
We chose to have the Evaluator rate on specific metrics because this strategy allows for iterative refinement of the prompt by adjusting the weight and threshold of each metric. This approach decreases uncertainty by introducing fixed numerical comparisons, which provides a clearer direction for tuning, as opposed to relying solely on textual standards, which can be more ambiguous.
In designing the Evaluator’s prompt, we started by selecting a small subset of human-labeled samples. We began with an initial weight distribution of ($0.7,0.2,0.1$) to reflect the relative importance of each metric in Section 2.4. The high initial weight for ”Precise Contextual Evidence” was chosen to align with our focus on whether the Curator identifies annotated issues by providing contextual evidence, as discussed in Section 2.4. We then fine-tuned these weights using the selected subset to better align the Evaluator’s prompt with human judgments.
The motivation behind this design is to make iterative refinement of prompts more effective. Developing a good prompt requires testing it on a subset of human-labeled issues with corresponding outcomes from Curators and then making adjustments based on the outcomes. Unlike the model training process, which benefits from clear gradients and systematic optimization, tuning text prompts is often an empirical process necessitating numerous trials without the guarantee of improvement in its performance. Employing numerical weights and thresholds could enable us to tune the prompt more fine-grained, only changing the weights or threshold based on the results, which share similarities with linear regression, resulting in a more efficient prompt tuning process and better performance. By ”training” on a this human-labeled subset, we finalized weights in the prompt as $0.85$, $0.15$, and $0.05$.
It’s important to highlight that we collected samples for the test set used in Section 3.1 after finalizing the prompt design, ensuring that there is no test data leakage to the prompt tuning.
To be more detailed, the Evaluator is prompted with the following instructions to evaluate the performance of the agent, with <ISSUE>, <HINT>, and <ANSWER> being the place-holder.
You are required to act as an answer evaluator. Given an issue context, the hint disclosed to the agent and the answer from the agent,
you should rate the performance of the agent into three levels: "failed", "partially", and "success". The rating rules are as follows:
<rules>
1. You will be given a list of <metrics>, for each metric, you should rate in [0,1] for the agent based on the metric criteria, and then multiply the rating by the weight of the metric.
2. If the sum of the ratings is less than 0.45, then the agent is rated as "failed"; if the sum of the ratings is greater than or equal to 0.45 and less than 0.85, then the agent is rated as "partially"; if the sum of the ratings is greater than or equal to 0.85, then the agent is rated as "success".
3. **<text>** means the text is important and should be paid attention to.
4. ****<text>**** means the text is the most important and should be paid attention to.
</rules>
The <metrics> are as follows:
"m1": {
"criteria": "Precise Contextual Evidence:
1. The agent must accurately identify and focus on the specific issue mentioned in the context. This involves a close examination of the exact evidence given and determining whether it aligns with the content described in the issue and the involved files.
2. Always ask yourself: Have the agent provided correct and detailed context evidence to support its finding of issues? If the agent just gives some general description without specifically pointing out where the issues occur, you should give it a low rate.
3. Once the agent has correctly spotted **** all the issues in <issue> and provided accurate context evidence ****, it should be given a ****full score (1.0) even if it includes other unrelated issues/examples ****"
4. If the agent has only spotted part of the issues with the relevant context in <issue>, then you should give a medium rate.
5. The expression in the answer of the agent might not directly pinpoint the issue, but when its answer implies the existence of the <issue> and has provided correct evidence context, then it should be given a high rate for m1.
6. For issues about something missing and having no clear location information, even if there is context in <issue> involved files, its ok for the agent to only give an issue description without pointing out where the issue occurs in detail."
"weight": 0.8,
"range": [0, 1]
"m2": {
"criteria": "Detailed Issue Analysis:
1. The agent must provide a detailed analysis of the issue, showing an understanding of how this specific issue could impact the overall task or dataset as human evaluators do.
2. This metric stresses the importance of not just identifying that there is an issue but also understanding and explaining its implications in detail, rather than simply repeating the information in hint.",
"weight": 0.15,
"range": [0, 1]
"m3": {
"criteria": "Relevance of Reasoning:
1. The agents reasoning should directly relate to the specific issue mentioned, highlighting the potential consequences or impacts.
2. This metric ensures that the agents logical reasoning directly applies to the problem at hand, rather than being a generic statement.",
"weight": 0.05,
"range": [0, 1]
</metrics>
Reference:
<issue>
<ISSUE>
</issue>
<tips-for-involved-field>
- In the <issue> "involved" field, the content of each file shows lines beginning with "+" or "-" to indicate lines that are newly added or removed to fix the issue. This means the file accessed by the agent is the previous version of the file.
- For issues that only involve adding lines, such issues often have a wider and more general context. Therefore, we should be less strict when evaluating m1, and give it a higher rate.
- For issues that only involve removing lines, such issues often have a narrower and more specific context. Therefore, we should be more strict when evaluating m1.
</tips-for-involved-field>
<hint>
<HINT>
</hint>
-------------------- Below is the answer from the agent. Ensure you dont take the information above as the agents answer!
<answer>
<ANSWER>
</answer>
response below,
1. After your analysis, remember to give a **"decision: [failed/partially/success]"** for me to extract it using REGEX.
2. Dont use Code Interpreter!; Use your ability to analyze the text. ** Pay attention to your calculations and make sure they are correct. **
3. There could be multiple issues described in <issue> part. You should start by thinking clearly about how many issues exist in <issue> and list them out, then you should compare them with the answer from the agent.
4. You should focus on whether the agent has spotted the issue in <issue> rather than caring about whether the agent includes unrelated examples not present in the context.
5. The agent is not tasked with solving the problem; its only asked to identify the issue. Therefore, its acceptable for the agent to merely describe the existence of the issue without providing a solution, as long as it is mentioned in the reference involved context.
Please note the categories of the Curator’s performance are different from papers. We use partially for success in the paper. This modification is made because we want to emphasize we focus more on separating fail from the other two categories, which is a binary classification problem. Using partially might cause confusion.
B.2. The Settings of Evaluators
When developing Evaluators, we test following different models as backends: GPT-4o-2024-1120, GPT-4o-2025-0513 (OpenAI et al., 2024), o3-mini-2025-01-31 (OpenAI, 2024b), GPT-4-0125 preview (OpenAI, 2024a), GPT-3.5-Turbo (OpenAI, 2023b), Llama3.3-70B-Instruct (AI, 2024), Llama3-70B-Instruct (Dubey et al., 2024) Deepseek R1 (DeepSeek-AI et al., 2025a), and Deepseek V3 (DeepSeek-AI et al., 2025b). For OpenAI series of models, we applied OpenAI official API [^6] https://platform.openai.com/docs/api-reference, and set their temperature to $0.0$ to maximize reproducibility and stability. For other open source models, we used SGLang framework (Zheng et al., 2024) for efficient deployment and inference. We didn’t apply qunatizations in order to keep model’s original performance. The dtype is Bfloat16, which is default value. We set temperature to $0.6$ for DeepSeek-R1, following their model cards [^7] https://huggingface.co/deepseek-ai/DeepSeek-R1, while setting temperature to $0.0$ for all other models.
B.3. The Settings of Curators
We implemented several baseline Curators using SGLang (Zheng et al., 2024), LangChain (Chase, 2022), and the OpenAI Assistant API. For open-source models, including DeepSeek-R1 (DeepSeek-AI et al., 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2025b), we deployed them locally using the SGLang framework.
To implement the Retrieval-Augmented Generation (RAG) pipeline, we utilized the LangChain framework. We first built a knowledge vector store using the FAISS (Douze et al., 2024) package and then constructed the retrieval chain using the create_stuff_documents_chain and create_retrieval_chain functions provided by LangChain. To integrate SGLang-deployed models within LangChain, we implemented a custom class named SGLangLLM, following LangChain’s model interface.
Regarding temperature settings, we set the temperature to $1.0$ for OpenAI models and DeepSeek-V3. For DeepSeek-R1, we followed its model card recommendations and set the temperature to $0.6$.
B.4. Human Annotation Codebook for Evaluating Curator’s Performance
The human annotators are required to follow the following codebook when rating the responses from the Curators:
You are tasked with evaluating the work of a Dataset Curator, who is responsible for identifying issues in dataset files and pinpointing the contextual evidence within those files. You will be provided with a ground truth that specifies the issues present in the files and their exact locations.
Based on this information, please categorize the Curator’s findings into the three following categories:
  • fail: the Curator either denies the existence of the issue stated on the label, identifies irrelevant issues, or acknowledges the issue but offers completely inaccurate supporting evidence.
  • success: The Curator identifies the annotated issue and provides at least one correct piece of contextual evidence.
  • success+: the Curator correctly identifies all issues and provides all necessary contextual evidence.
Additional Rules: 1. For issues about something missing and having no clear location information, even if there is context in ¡issue¿ involved files, it’s ok for the agent to only give an issue description without pointing out where the issue occurs in detail. 2. You should focus on whether the agent has spotted the issue in ground truth rather than caring about whether the agent includes unrelated examples not present in the context. The inclusion of other irrelevant issues won’t influence its performance in evaluation.
B.5. Additional Results from Human-Alignment Experiment on Evaluator
As defined in Section 2.3, the performance of the Curator is categorized into fail, success, and strict-success. However, accurately distinguishing between success and success+ in the original triple classification is challenging for Evaluators.
Tab. 6 presents the complete results of the Evaluators’ performance with success+ alignment, compared to human annotations (ground truth). There is a significant drop in the Evaluators’ performance when required to provide a triple classification.
Table 6. The performance of Evaluators with different backend models (Binary and Triple Classification), using human annotations as ground truth. We treat succ as the positive class when calculating Precision, Recall, and F1 Score (Binary). Color Legend:   (95–100%),   (85–95%),   (60–85%),   (0–60%). Compared with binary classification (fail / (success+success+)), the triple classification proves to be more difficult, making the evaluation results from the Evaluator less reliable, we therefore leave that part for future works.
Binary / % Triple / %
Model Name Accuracy Precision Recall F1 Score $\kappa$ Value Accuracy $\kappa$ Value
gpt-4o-2024-11-20 97.83 100.00 92.59 96.15 94.64 88.04 72.98
gpt-4-0125-preview 96.74 92.86 96.30 94.55 92.22 88.04 73.77
gpt-4o-2024-05-13 92.39 81.25 96.30 88.14 82.59 83.70 66.46
DeepSeek-R1 92.39 81.25 96.30 88.14 82.59 86.96 72.61
DeepSeek-V3 93.48 88.89 88.89 88.89 84.27 83.70 63.99
gpt-3.5-turbo 68.48 48.21 100.00 65.06 42.15 63.04 38.82
Meta-Llama-3-70B-Instruct 69.57 49.09 100.00 65.85 43.68 64.13 41.30
Meta-Llama-3.3-70B-Instruct 88.04 72.22 96.30 82.54 73.73 80.43 61.52
o3-mini-2025-01-31 91.30 100.00 70.37 82.61 77.04 82.61 56.96
B.6. Consistent Evaluation
To evaluate the consistency of ratings produced by this prompting strategy, we ran the Evaluator, backed by GPT-4-0125-preview and GPT-4o-2024-11-20, separately three times on the same outputs generated by the Baseline Curator (GPT-4-0125-preview) using a subset of DCA-Bench (92 data-hint combinations). For GPT-4-0125-preview, the standard deviation is ±$\pm 0.89\%$, while for GPT-4o-2024-11-20, the standard deviation is ±$\pm 0.51\%$, indicating high stability. The detailed numbers are included in Tab. 7.
Table 7. Success rates and standard deviations of multiple runs with different backend models (in percentage)
Model Trial 1 (%) Trial 2 (%) Trial 3 (%) Std Dev (%)
GPT-4-0125-preview 27.17 26.09 28.26 0.89
GPT-4o-2024-1120 27.17 26.09 26.09 0.51
B.7. Negligible Self-Preference and Length Bias of Evaluator
In this section, we discuss the possibility that LLM Evaluators may be biased towards Curators that are backed by the same LLM model (self-preference), and might be biases toward the length of the response (length bias) (Panickssery et al., 2024; et al., 2023). We design two experiments to examine whether these biases are within acceptable range.
Experiment on Self-Preference Bias We used GPT-3.5-Turbo and GPT-4-0125-preview as the backend models for both the Curators and the Evaluators, leading to four combinations of Curator and Evaluator settings. For each Evaluator, we calculated the Average Odds Difference (AOD) (Majumder et al., 2022), a metric that measures fairness by assessing differences in true positive and false positive rates between two groups. A value of zero indicates perfect fairness, with equal model performance across both groups. Higher AOD values suggest increasing disparity, with the model performing unevenly between the groups. The AOD is defined as:
(1) $\text{AOD}=\frac{1}{2}\left(\left|\text{TPR}_{\text{group1}}-\text{TPR}_{\text{group2}}\right|+\left|\text{FPR}_{\text{group1}}-\text{FPR}_{\text{group2}}\right|\right)$
In this context, the GPT-4-0125-preview Curator is privileged, while the GPT-3.5-Turbo Curator is unprivileged. We evaluated the GPT-3.5-Turbo Evaluator five times and the GPT-4-0125-preview Evaluator three times.
Table 8. True Positive Rate (TPR) Results
Evaluator \ Curator GPT-3.5-Turbo GPT4-0125-preview
GPT-3.5-Turbo $100.00\,(\pm 0.00)\%$ $100.00\,(\pm 0.00)\%$
GPT-4-0125-preview $84.44\,(\pm 3.14)\%$ $91.35\,(\pm 3.49)\%$
Table 9. False Positive Rate (FPR) Results
Evaluator \ Curator GPT-3.5-Turbo GPT4-0125-preview
GPT-3.5-Turbo $68.95\,(\pm 1.29)\%$ $56.78\,(\pm 3.18)\%$
GPT-4-0125-preview $0.00\,(\pm 0.00)\%$ $3.71\,(\pm 2.92)\%$
Using the error propagation formula [^8] https://en.wikipedia.org/wiki/Propagation_of_uncertainty and Eq. 1, we get the AOD for GPT-3.5-Turbo is $6.09\,(\pm 1.72)\%$.We can also get that the AOD for GPT-4-0125-preview, which is $5.31(\pm 2.76)\%$.
Since an $AOD\leq 10\%$ is considered fair (Majumder et al., 2022), the bias remains within an acceptable range.
Furthermore, we observe a great alignment between the GPT-4-0125-preview based Evaluator and human annotations, on both the GPT-3.5-Turbo based Curator and the GPT-4-0125-preview based Curator:
Table 10. Comparison of Metrics between Evaluators and Human Annotations
Metric/Curator GPT-3.5-Turbo GPT-4-0125-preview
Accuracy $97.46\,(\pm 0.51)\%$ $96.38\,(\pm 0.51)\%$
Precision $100.00\,(\pm 0.00)\%$ $96.29\,(\pm 2.92)\%$
Recall $84.44\,(\pm 3.14)\%$ $91.36\,(\pm 3.49)\%$
F1 Score $91.53\,(\pm 1.87)\%$ $93.66\,(\pm 0.97)\%$
$\kappa$ value $90.06\,(\pm 2.16)\%$ $91.13\,(\pm 1.31)\%$
These results suggest that using the same model for dataset curation and evaluation does not introduce significant bias in our study, which is acceptable.
Experiment on Length Bias We also conducted an experiment to further investigate this. After the Curator generated the results, we fed its response to another LLM (gpt-4o-2024-08-06) to expand upon its generation. While the most common bias is the preference towards longer responses, we also tested the case where the response is more succinct. The corresponding prompts are listed as follows.
Rephrase to verbose:
You are provided with a description text, and your task is to make it more detailed and elaborate. Ensure that you do not add any new information or content that was not present in the original context. Your goal is to expand upon the existing meaning without altering it. Additionally, maintain the same format as the original answer. Dont mention this prompt in your response.
=================== Start:
Rephrase to succinct:
You are provided with a description text. Your task is to rephrase it concisely while ensuring all essential information is retained. Do not add new information or omit any key details from the original context. Remember to maintain and keep the same format as the original answer. When answer is already short, you dont have to condense it too much. Do not mention this prompt in your response.
=================== Start:
Both prompts are designed to ensure that the only difference between the responses is the length. We used GPT-3.5-Turbo (0125) and GPT-4o-mini (0718) as the underlying models for the Curator and compared the results with responses of different length to see if there was a significant change in the success rate. The experiment was conducted on 92 cases (at the hint level).
Table 11. Average response length of succinct, normal, verbose settings.
Model Name/Running Type Succinct Normal Verbose
GPT-3.5-Turbo $204.8$ $339.1$ $520.0$
GPT-4o-mini $361.4$ $521.8$ $672.9$
Table 12. Accuracy of succinct, normal, verbose settings.
Model Name/Running Type Succinct Normal Verbose
GPT-3.5-Turbo $13.04\%$ $14.13\%$ $13.04\%$
GPT-4o-mini $20.65\%$ $20.65\%$ $19.57\%$
As can be seen, the difference of accuracy is at most $\approx 1\%$ for various response lengths, indicating that there is no significant length bias with our Evaluator.
B.8. Prompt Design of Hints Writer
The Hints Writer is prompted with the following instructions to generate three level hints for the issue, with <ISSUE> being the placeholder.
You are an issue-hint writer. You need to formulate some hints for detecting certain issues, based on the description in
<issue> below.
<rules>
- The hints should have three different levels, from very general to very specific.
- For ‘‘general hint’’, you should only tell the general issue, without any specific information.
- For ‘‘medium hint’’, you could tell which files are involved in this issue + general hint. Make sure it contains the information from the general hint.
- For ‘‘specific hint’’, you should tell which files are involved in this issue clearly, as well as a part of context information. Make sure it contains the information from the general and medium hints.
- You should never tell the agent the full context of the issue, only give it hints, so we can test to what degree the agent detects the issue and tell us evidence by themselves.
- Your word used in hints should be precise, pay attention to the word choice, and try to avoid misleading or ambiguous words. You should make sure you have a clear understanding of the issue before writing the hints.
- The hints should be clear and concise, try to cover all the issues mentioned and avoid unnecessary information.
</rules>
<example>
<issue>
title: What is the provenance of these benchmark datasets?}
content:
Firstly, check the English proverbs dataset: https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/english_proverbs
task can help to draw conclusions about human-like understanding and
## Data source
## References
There is nothing in the data source section.
involved: {
name: README.md,
context: {
task can help to draw conclusions about human-like understanding and Data Source References
</issue>
output:
general: "section with empty content"
medium: "a section with empty content in markdown file"
specific: "a sub-heading section in README.md with empty content"
</example>
Now lets begin
<issue>
<ISSUE>
</issue>
write your hints here in JSON format:
When designing the prompt, we take into consideration the risk that hints could provide too much information that can already serve as a valid answer. In the “¡rule¿” section, we clearly require the Hint Writer not to reveal the full context of the issue to the Curator, in case it takes it as a shortcut and skips reading the real files. We also manually check the generated hints to ensure the generated hints comply this rule.
Appendix C Detailed Analysis of the Performance of Baseline Curator on DCA-Bench
C.1. From the Perspective of Different Content Length
Given the large average content length (or size of files, counted by token number) involved in test cases, it’s interesting to explore how the success rate of the Baseline Curator changes with the number of tokens. Therefore, we created a statistical plot showing the count of success and failure cases for different content lengths. Fig. 3 shows that the success rate doesn’t monotonically decrease as content length grows. For our baseline Curators, we used the OpenAI Assistant API with the Code Interpreter Tool to build the agents. This tool provides temporary file storage for uploaded files and can process them with code scripts. By processing data before it is passed to the LLM, the tool allows the OpenAI Assistant to handle much smaller inputs in terms of token usage, even when working with large datasets. Consequently, the effective input size for the LLM is often significantly reduced. Notably, depending on the agent’s capacity to efficiently process and manage information.
Figure 3. Number of success & failure and success rate versus different content length of test cases.
Besides, the performance of the Curator also depends on the type of issues. Take the issue 21ca944c-cf82-4764-bb2c-4c8db0cee950999https://www.kaggle.com/datasets/roche-data-science-coalition/uncover/discussion/168946 as an example. This issue involves many missing values in a large CSV file. Even with 361,814 tokens, it’s relatively easy for an LLM agent to spot the problem. The agent can quickly check part of the file and identify the missing values. This example shows that task difficulty isn’t just about data size. It also depends on the problem type and how relevant the information is.
Therefore, it’s expected that a test case with larger content length does not necessarily lead to poorer performance.
C.2. From the Perspective of Hint Level
Here is a more detailed analysis of the overall performance with different hints. As outlined in L183-L186, the hints are divided into four levels, from 0 to 3, focusing on two dimensions with increasing information: the description of the issue and its specific context. The design follows an incremental pattern from h0 to h3. Please refer to Appendix 10 for some specific examples. Let’s continue by discussing the capabilities required to succeed at each hint level.
If we aim to develop an LLM agent for this task without considering cost constraints, the process should involve the following steps: Retrieve Knowledge → Plan and Select Files to Check → Read Files and Validate Issues → Receive Feedback and Update State → (Loop) → Make Final Decisions. Since our Baseline Curator does not support looping, we exclude it from the following analysis. We will use this diagram to demonstrate which steps become less critical as the hint level increases.
h0: With no hints provided, the Curator must examine all the uploaded files. This requires the ability to handle and retain very long inputs effectively. To identify potential issues, the Curator needs comprehensive knowledge of dataset curation to anticipate a wide range of problems and retrieve relevant information when provided with the dataset files. The task also involves exploring and interacting with the dataset files to gradually narrow down the search space, requiring high-level reasoning and planning abilities. This corresponds to no reduction in the original diagram’s required steps.
h1: When given a general description of the issues but no specific location, the search space is somewhat reduced but not entirely, as the description remains broad. A robust understanding of dataset curation and the ability to manage long inputs are still necessary. Ideally, this follows the pattern of Retrieval of Knowledge ($\downarrow$), with minimal impact on the steps for planning and selecting files, except where file names are relevant. The steps for reading files and validating issues also see a slight reduction ($\downarrow$).
h2: At this level, the Curator is provided with both a general description of the issue and specific hints about which files are involved. This further reduces the search space, making the ability to manage long inputs sufficient but less critical. The need for extensive knowledge retrieval decreases as the Curator can focus on the relevant files. Planning and file selection are simplified, as the Curator has clear guidance on which files to examine. Reading and validating issues also become easier, as the scope is narrower, reducing the overall complexity of the task. This follows the pattern of Retrieval of Knowledge ($\downarrow$), Planning and Choosing Files ($\downarrow$), and Reading and Validating Issues ($\downarrow$).
h3: At the highest hint level, the Curator receives the general description, the specific files involved, and partial context. This greatly reduces the need for managing long inputs and makes the retrieval of knowledge more targeted. Planning and file selection is almost straightforward, with the Curator essentially guided to the areas needing attention. Reading and validating issues are significantly simplified due to the detailed context, further decreasing the Curator’s workload. This follows the pattern of Retrieval of Knowledge ($\downarrow$ $\downarrow$), Planning and Choosing Files ($\downarrow$ $\downarrow$), and Reading and Validating Issues ($\downarrow$ $\downarrow$).
C.3. From the Perspective of Different Issue Tags
We show the evaluation results from the Baseline Curator on the whole DCA-Bench classified into different tags in Tab. 13.
Table 13. The success rate of Curator on DCA-Bench (221 samples $\times$ 4 hint levels) grouped by tags. The Curator is backended by GPT-4-0125-preview with OpenAI Assisstant API. The Evaluator is backended GPT-4-0125-preview.
Category Success Rate/Number Sub-category Success Rate/Number
typo 33.33% / 18*4
wrong-format 46.43% / 14*4
inappropriate-file 50.00% / 4*4
ethical/legal-risk 47.50% / 10*4
internal-discrepancy 36.90% / 21*4
cross-file-discrepancy 28.41% / 44*4
data-problem 32.99% / 197*4 wrong-value 29.58% / 71*4
missing-value 40.00% / 15*4
data-leakage 50.00% / 2*4
apparent-corruption 45.63% / 40*4
hidden-corruption 25.42% / 59*4
document-problem 40.96% / 83*4 wrong-info 29.63% / 27*4
insufficient-info 45.67% / 52*4
infrastructure-problem 35.53% / 19*4 data-access 56.25% / 4*4
script-code 30.00% / 15*4
We conduct analysis for some tags with relatively low success rates below.
Typo This refers to issues caused by typographical errors, including misspelled words, misused phrases, and incorrect email addresses. The specific nature of these issues is that they involve small errors within a long context, which means a large search space. This makes errors that are easy to spot in shorter contexts extremely difficult to find, even for humans.
Internal-discrepancy This involves issues within a file where information is misaligned, such as a mismatched sum for some features in a .csv file [^10] https://github.com/fivethirtyeight/data/issues/52, or a word misused in a way that contradicts itself within its context [^11] https://github.com/google/BIG-bench/pull/904/files#diff-a5582cf9c7731f66cc3e8f6aa0e9d070483fc14f7c1d3545f275207c6fa39bd5. These issues require reasoning and memory over a long context.
Cross-file-discrepancy Compared to internal-discrepancy, this type of issue requires checking for consistency across different files, such as misaligned information between a datacard and dataset files [^12] https://github.com/fivethirtyeight/data/issues/261. This can be challenging when there are many issues and limited hints about the files involved.
Wrong-info This involves incorrect information in documentation files (e.g., wrong meta-information, typos, invalid URLs, and email addresses). This category overlaps with the issues mentioned above.
Script-code Many datasets also provide script code to load the data, and issues can arise in these scripts, leading to data loading problems [^13] https://huggingface.co/datasets/spyysalo/bc2gm_corpus/discussions/3. To address these issues, the Curator needs to be skilled in coding and debugging and sometimes must understand the dataset’s purpose [^14] https://github.com/google/BIG-bench/pull/559.
Appendix D Ethics Statement
D.1. Research involving human participants
The human annotators are the authors of this work, and we have confirmed that they are willing to annotate, knowing the potential ethically problematic data issues, as we listed in Offensive Content. Prior to the annotation, the leading authors have checked that the datasets with ethical problems pose minimal risks to the annotators.
D.2. Data Privacy, Copyright, and Consent
The DCA-Bench has collected materials and datasets from the following platforms: HuggingFace, Tensorflow Dataset, Graph Learning Indexer, Kaggle, BIG-bench, OpenML, Open-Data-Registry, and FiveThirtyEight. The corresponding terms or license links are provided.
Besides, we have also included a detailed table DCA_Benchmark.csv in our supplementary materials, where all the files involved in DCA-Bench have been annotated with their License Information.
Each data point in DCA-Bench has two types of license:
  • License for the platform that hosted this dataset
  • License of this dataset
Details:
  • Some data points in DCA-Bench involve files from 2 datasets, we listed all licenses.
  • Many datasets themselves don’t have a license listed on their data card.
  • We notice that there are some datasets151515e.g. https://www.kaggle.com/datasets/roche-data-science-coalition/uncover/dataclaiming that ”Data files © Original Authors”, which is not a standard License. We have reached out to the dataset owners for clarification, but have not yet received a response. Therefore, we choose to record them as it claimed in our table.
How does this secondary usage of user-generated data comply with restrictions?  DCA-Bench involves user-generated data (comments, modified codes) collected from dataset repositories hosted on GitHub, HuggingFace, and Kaggle.
For GitHub, we collected the comments and modified codes generated by users. According to section D.3, paragraph 2 of GitHub Terms of Service [^16] https://docs.github.com/en/site-policy/github-terms/github-terms-of-service#c-acceptable-use,
”Because you retain ownership of and responsibility for Your Content, we need you to grant us—and other GitHub Users—certain legal permissions, listed in Sections D.4—D.7.”
According to section D.5. License Grant to Other Users [^17] https://docs.github.com/en/site-policy/github-terms/github-terms-of-service#5-license-grant-to-other-users, if not provided a specific license, any User-Generated Content you post publicly, including issues, comments, and contributions to other Users’ repositories, may be viewed by others. However, it doesn’t clearly explain how this content is allowed to be used in which ways, and which is not. However, according to a Github issue discussion [^18] https://github.com/orgs/community/discussions/135466, we believe our practice is acceptable. Besides, we noticed that there have already been some works (Kocetkov et al., 2022; Jimenez et al., 2024b) with similar usage of GitHub data, which implies that this usage is acceptable.
For HuggingFace, according to HuggingFace Content Policy [^19] https://huggingface.co/content-guidelines, Content types may include:
  • ”ML Artifacts”: Code and assets hosted as Hugging Face Repositories, including Models, Datasets, and Spaces;
  • ”Community Content”: Content that can be found in the Community section of the Hugging Face Platform, including discussions, comments, and usernames, as well as related documentation such as READMEs, model cards, data cards, pull requests, and merges.
According to HuggingFace Terms of Service [^20] https://huggingface.co/terms-of-service, Section “Your Content”,
”If you decide to set your Repository public, you grant each User a perpetual, irrevocable, worldwide, royalty-free, non-exclusive license to use, display, publish, reproduce, distribute, and make derivative works of your Content through our Services and functionalities;”
Therefore, we believe our usage of user-generated content from HuggingFace should belong to the “derivative works” here, which is acceptable.
For Kaggle, after reviewing their Terms of Use [^21] https://www.kaggle.com/terms, we noticed that there are no explicit guidelines regarding the use of user-submitted content for academic research purposes. So we have sent Kaggle Support a request for clarification, and here is the response:
” Kaggle is a neutral platform in hosting competitions and data. Protections under copyright law can vary based on the nature of data, contractual obligations, and your jurisdiction. As a result, we are not able to approve individual uses of data on a case-by-case basis. Copyright or usage questions should be directed to your legal counsel. You may post usage enquiries in the respective competition’s forum, but we can not guarantee an official response from the Sponsor.
If the rules do not specify usage restrictions, we ask that you use your best judgement in determining fair use. For example, it may be okay for you to use the data for teaching a class, demoing a method in a blog post, or other non-commercial purposes. Some of our competitions identify themselves as open-source and are meant to foster research, collaboration, and continued analysis by the data science community. ”
Since our paper is not for commercial usage, we believe our practice is acceptable
Lastly, our collection and evaluation processes exclude the gathering of any user information. We promise that we will remove the content once requested with a valid reason.
D.3. Offensive Content
Our dataset, comprising 221 data instances, includes content that may be considered sensitive or potentially controversial. Specifically, there are 10 data points involving ethical or legal risks, among which 2 data points222222 https://github.com/google/BIG-bench/pull/685 and https://www.kaggle.com/datasets/vikrishnan/boston-house-prices/discussion/429030 contain content that may exhibit bias towards specific groups of people. Other issues are relevant to data source or license information.
Besides, the risk in most cases is associated with deploying models trained with these datasets. Reviewing the quality issues of these datasets poses minimal risks to the annotators.
Justification for Inclusion
While we acknowledge the sensitive nature of these data instances, their inclusion in our dataset is both intentional and necessary for the following reasons:
  • Benchmark Objectives: One of the primary goals of our benchmark is to examine LLM agents’ ability to identify and assess potential ethical and legal risks in AI training data. The inclusion of these sensitive data points is crucial for thoroughly evaluating the capability of AI models to recognize and appropriately handle such content.
  • Realistic Representation: These data points reflect real-world scenarios that AI systems may encounter. By including them, we ensure our benchmark provides a more comprehensive and authentic assessment of AI performance.
Appendix E Other Techinal Details
E.1. Discussion on the Difference between OpenAI Assistant API and ChatGPT
The performance discrepancy between OpenAI’s Assistant API and ChatGPT has been widely discussed within the OpenAI user community. Users often report that the Assistant API’s performance does not match the capabilities of ChatGPT, especially in tasks involving complex file processing and analysis.
The Assistant API provides several methods for handling files: 1. Code Interpreter for file uploads. 2. Building VectorStores via File Search. 3. Attaching files to messages using their file IDs.
However, these methods have significant issues: - The quality of the VectorStore built using File Search is inferior to ChatGPT’s built-in file search capabilities. - Files are referenced by their file IDs rather than their names, leading to confusion and errors during processing.
The following code snippet illustrates the process of uploading and referencing files in the Assistant API by attaching them to the messages:
msg_file_ids = []
for file_path in file_paths:
message_file = self.client.files.create(
file=open(file_path, "rb"), purpose="assistants"
msg_file_ids.append(message_file.id)
attachments = [{"id": file_id, "tools": [{"type": "file_search"}]} for file_id in msg_file_ids]
thread = self.client.beta.threads.create(
messages=[
"role": "user",
"content": input,
"attachments": attachments
Despite correctly assigning file names during the upload process, the Assistant API seems to ignore these names during subsequent file processing, relying solely on file IDs. This results in frequent mix-ups and incorrect file references.
Several forum posts and articles highlight deficiencies relevant to this topic: - Different responses assistant playground vs API - using the same assistant ID [^23] https://community.openai.com/t/different-responses-assistant-playground-vs-api-using-same-assistant-id/547562 - Huge difference between ChatGPT assistant and API assistant [^24] https://community.openai.com/t/huge-difference-between-chatgpt-assistant-and-api-assistant/587335 - Assistant API not consistent [^25] https://community.openai.com/t/assistant-api-not-consistant/718145. This implies the limitations of OpenAI Assistant API.
In conclusion, while the Assistant API offers various tools for convenient file handling, its current implementation presents significant challenges that hinder its effectiveness compared to ChatGPT, which motivates us to select a subset of DCA-Bench containing hard cases to test on ChatGPT.
E.2. Calculation of File Number and Context Length
We calculate the file context length using the tiktoken package [^26] https://github.com/openai/tiktoken to tokenize the file content, using:
import tiktoken
def calculate_token_len(sentence, model="gpt-4-0125-preview"):
encoding = tiktoken.encoding_for_model(model)
tokens = encoding.encode(sentence)
token_len = len(tokens)
return token_len
During calculation, non-text files are skipped, which include .jpg, .jpeg, .png, .gif, .bmp, .mp3, .wav, .ogg, .flac, .mp4, .avi, .mov, .mkv.
For .zip files, we will unzip them to calculate the context length of text files. However, when calculating the file number, we take the whole .zip number as one single file.
E.3. Cost Analysis
In this section, we provide an estimated cost analysis for applying our Evaluator backend by GPT-4-0125-preview to rate the performance of the Baseline Curator (OpenAI Assistant API) on DCA-Bench (884 inputs in total). We applied tiktoken for tokenization of contexts.
Based on the pricing information available as of May 2024 from https://openai.com/api/pricing/, the costs are: input cost: \$16.12, output cost: \$28.99, total cost: \$45.11.
Using the same calculation methods, we obtain the following costs for GPT-4o-2024-1120: Input cost: \$4.46, Output cost: \$15.60, Total cost: \$20.06.
The costs are rounded to two decimal places, and the actual cost may vary depending on the output context generated by the Curator. Additionally, since OpenAI now supports cached inputs, the actual cost may be lower than the calculated value when the Curator’s outputs remain unchanged. It is important to note that the cost estimates provided here pertain only to the Evaluator; we exclude the cost of running the baseline Curator on DCA-Bench to obtain the outputs, as this cost depends on the Curator’s design and is not the focus of our analysis.